OpenStudy (anonymous):

Solving radical equations... Solve each equation: 1/4 - x = x/8 2X/3 - 1/2 = 2X+5/6

OpenStudy (anonymous):

I meant rational

OpenStudy (p0sitr0n):

x =2/9 x =-1

OpenStudy (anonymous):

thank you but how do I do it?

OpenStudy (p0sitr0n):

1/4 = x/8 + x 1/4 = x/8 + 8x/8 1/4 = 9x/8 1 = 9x/2 2 = 9x 2/9 =x

OpenStudy (p0sitr0n):

2x/3-2x = 1/2 +5/6 2x/3 -6x/3 = 3/6 +5/6 -4x/3 = 8/6 -4x = 4 x =-1